In a town election with two candidates, the winning candidate had 220 more votes than the loser. If 810 votes were case in all, how many votes did each candidate receive?
Found 2 solutions by 1980.vikas, Bucket:
Answer by 1980.vikas(7)

Let the loser get x no. of votes, then the winner gets (x + 220) votes.
Now as the total no of votes is 810, then
=> x + (x + 220) = 810
=> 2x + 220 = 810
=> 2x = 590
=> x = 295
so the loser get 295 votes while the winner gets 515 votes.

Answer by Bucket(1)

(810-220)/2 = 295 Cand1
295+220 = 515 Cand2